# Picogram-Level Nanoplastic Analysis with Nanoelectromechanical System Fourier Transform Infrared Spectroscopy: NEMS-FTIR

### v2

@@ Acknowledgements
The authors sincerely thank Jürgen Smoliner, Alois Lugstein and Masiar Sistani from the Institute of Solid State Electronics at TU Wien for letting us use their FTIR spectrometer. The authors gratefully acknowledge Noelia Barrabés Rabanal and the Research Group for Model Catalysis and Applied Catalysis at TU Wien for kindly providing access to their ATR module and supporting our measurements. The authors thank Michael Schneider and Paolo Martini from the Institute of Sensor and Actuator Systems at TU Wien for acquiring the SEM images and helping with finite element method simulations, respectively. Special thanks go to the Invisible-Light Labs team for their technical expertise and supportive encouragement, as well as to colleagues from the MNS group at TU Wien (Institute of Sensor and Actuator Systems) for their valuable insights and collaboration. The authors also gratefully acknowledge Georg Pfusterschmied for providing the image in Figure 1A. The authors gratefully thank Mihnea Surdu from the Extreme Environments Research Laboratory at EPFL in Switzerland for his assistance with the aerosol sampling setup.
Funding: This project received funding from the European Commission under grant agreement HORIZON–EIC–2021–TRANSITIONOPEN–101058011 – NEMILIES and 101202787 EMILIE Booster, the Austria Wirtschaftsservice (AWS SEED, Deep Tech & Physical Sciences project P2422873), the Austrian Science Fund (FWF) under project I 6086–N, and the Novo Nordisk Foundation under grant number NNF22OC0077964 – MASMONADE.
@@
Writing-review & editing: JTP, JPL, NL, and SS.
Competing interests: NL, HB, JPL, and SS are co–founders of Invisible–Light Labs GmbH who provided the analytical instrumentation used in this study. JTP, NL, and JH are employed by Invisible–Light Labs GmbH. EŠ and AG declare no competing interests.
@@
Jelena Timarac–Popović1,2*, Johannes Hiesberger1,2, Eldira Šesto1, Niklas Luhmann2,
Ariane Giesriegl1, Hajrudin Bešić1, Josiane P. Lafleur2, and Silvan Schmid1‡‡‡Correspondence e–mail address: jelena.popovic@tuwien.ac.at, silvan.schmid@tuwien.ac.at
@@
(Dated: July 19, 2025)
A. NEMS–FTIR spectra of single NPs
The vertical lines observed in the NEMS–FTIR spectra (visible on Figure 4C) correspond to characteristic peak positions (in wavenumbers) associated with specific nanoplastic particles. Some lines correspond to potential minor oxidation or contamination during sample preparation.
The NEMS–FTIR spectrum of PS nanoparticles exhibit vertical lines at the following wavenumbers: 700, 753, 1452, 1493, 1601, 1745, 1800, 1868, 1943, 2849, 2923, 3026, 3060, and 3081 .
The NEMS–FTIR spectrum of PP nanoparticles exhibit vertical lines at the following wavenumbers: 1115, 1220, 1254, 1377, 1459, 1565, 1726, 2839, 2868, 2920, and 2952 .
The NEMS–FTIR spectrum of PVC nanoparticles exhibit vertical lines at the following wavenumbers: 1100, 1254, 1334, 1427, 1564, 2849, 2910, and 2971 .
B. NEMS–FTIR spectra of tea bag leachates
The NEMS–FTIR spectra of tea bag leachates (see Figure 5C) exhibit vertical lines at the following wavenumbers: 575, 990, 1116, 1170, 1210, 1235, 1282, 1371, 1437, 1460, 1553, 1642, 1720, 2860, 2930, 3087, and 3300 .
C. Derivation of the decadic linear attenuation coefficient for sparse particle film

### v3

@@ Acknowledgements
We sincerely thank J. Smoliner, A. Lugstein and M. Sistani from the Institute of Solid State Electronics at TU Wien for letting us use their FTIR spectrometer. We gratefully acknowledge N. Barrabés Rabanal and the Research Group for Model Catalysis and Applied Catalysis at TU Wien for kindly providing access to their ATR module and supporting our measurements. We thank M. Schneider and P. Martini from the Institute of Sensor and Actuator Systems at TU Wien for acquiring the SEM images and helping with finite element method simulations, respectively. Special thanks go to the Invisible-Light Labs team for their technical expertise and supportive encouragement, as well as to our colleagues from the MNS group at TU Wien (Institute of Sensor and Actuator Systems) for their valuable insights and collaboration. We also gratefully acknowledge G. Pfusterschmied for providing the image in Fig. 1B. We gratefully thank M. Surdu from the Extreme Environments Research Laboratory at EPFL in Switzerland for his assistance with the aerosol sampling setup.
Funding: This project received funding from the European Commission under grant agreement HORIZON-EIC-2021-TRANSITION-OPEN-101058011-NEMILIES and EIC-Transition-Booster-EMILIE-101202787, the Austria Wirtschaftsservice (AWS SEED, Deep Tech & Physical Sciences project P2422873), the Austrian Science Fund (FWF) under project I 6086-N, and the Novo Nordisk Foundation under grant number NNF22OC0077964 - MASMONADE.
@@
Writing-review & editing: JTP, JPL, and SS.
Competing interests: NL, HB, JPL, and SS are co-founders of Invisible-Light Labs GmbH who provided the analytical instrumentation used in this study. JTP, NL, HB, and JH are employed by Invisible-Light Labs GmbH. EŠ and AG declare no competing interests.
@@
Jelena Timarac–Popović1,2*, Johannes Hiesberger2, Eldira Šesto1, Niklas Luhmann2,
Ariane Giesriegl1, Hajrudin Bešić1,2, Josiane P. Lafleur2, and Silvan Schmid1*††footnotetext: Correspondence e–mail address: jelena.popovic@tuwien.ac.at, silvan.schmid@tuwien.ac.at
@@
(Dated: January 21, 2026)
A. Derivation of the decadic linear attenuation coefficient for sparse particle film
@@
| Type of plastics | Peak position (cm-1) | Assignment | References |
| PS | 700 | Aromatic ring bending | [76, 75, 77, 74] |
| 753 | Aromatic out-of-plane C–H bending | [76, 75, 77, 74] | |
| 1028 | C–H in-plane bending of the phenyl ring | [75, 77, 78] | |
| 1452 | C–H bending (scissoring, aliphatic backbone) | [75, 77, 78, 74] | |
| 1493 | Aromatic C–C stretching (ring vibration) | [76, 75, 77, 78, 74] | |
| 1583 | [75, 77] | ||
| 1601 | [76, 75, 77, 78, 74] | ||
| 1745 | Aromatic overtone/combination band (benzene fingers) | [76, 77] | |
| 1800 | [76, 75, 77] | ||
| 1868 | [76, 75, 77] | ||
| 1943 | [76, 75, 77] | ||
| 2849 | Aliphatic C–H stretching (CH2 asymmetric and symmetric) | [76, 75, 77, 74] | |
| 2923 | [76, 75, 77, 74] | ||
| 3026 | Aromatic C–H stretching | [76, 75, 77, 74] | |
| 3060 | [76, 75, 77] | ||
| 3081 | [76, 75, 77] | ||
| PP | 1115 | C–C chain stretching + CH3 rocking + CH2 wagging + CH twisting + CH bending | [79] |
| 1167 | C–H bending + C–C chain stretching + CH3 rocking | [84, 78, 79] | |
| 1220 | CH2 twisting + CH bending + C–C chain stretching | [79] | |
| 1254 | CH bending + CH2 twisting + CH3 rocking | [79] | |
| 1377 | CH3 bending, symmetric (umbrella mode) | [84, 78, 80] | |
| 1459 | CH3 bending, symmetric | [84, 78, 80] | |
| 2839 | CH2 stretching, symmetric | [76, 84] | |
| 2868 | CH3 stretching, symmetric | [76, 84, 80] | |
| 2920 | CH2 stretching, asymmetric | [76, 84, 80] | |
| 2952 | CH3 stretching, asymmetric | [76, 84, 80] | |
| PVC | 614 | C-Cl stretching | [81, 78] |
| 637 | [81, 78] | ||
| 692 | [81, 83, 78] | ||
| 970 | CH2 rocking | [81, 83, 84, 78] | |
| 1100 | C-C stretching | [81, 83, 84] | |
| 1254 | C-H bending, in-phase | [81, 83, 84] | |
| 1334 | C-H bending, out-of-phase | [81, 82, 83, 84] | |
| 1427 | CH2 bending, in-phase | [81, 83, 84] | |
| 1625 | C=C stretching | [82] | |
| 2849 | CH2 stretching, symmetrical, in-phase | [83] | |
| 2910 | C-H stretching of CH2, asymmetrical, in-phase | [81, 83, 84] | |
| 2971 | C-H stretching of CHCl, out-of-phase | [81, 83] | |
| PA | 575 | C=O out-of-plane vibration (amide VI) | [87, 88] |
| 990 | C-CO stretching (amide IV) | [87, 88, 89] | |
| 1116 | C-N stretching | [90, 86] | |
| 1170 | C-C stretching | [87] | |
| 1210 | CH2 wagging or twisting | [87] | |
| 1235 | C-N stretching | [86] | |
| 1282 | C-N stretching + CH2 twisting | [86, 87] | |
| 1371 | CH2 wagging + C-N-H deformation | [87, 89, 88] | |
| 1437 | CH2 (Gauche bonds) deformation | [87] | |
| 1460 | CH2 (adjacent to N-H group) deformation | [87, 88] | |
| 1553 | N-H in-plane bending + C-N stretching (amide II) | [86, 87, 88] | |
| 1642 | C=O stretching (amide I) | [86, 87, 88] | |
| 1720 | Imide group; carboxylic acid C=O stretching mode | [89, 91, 92] | |
| 2860 | C-H stretching, symmetrical | [87, 88] | |
| 2930 | C-H stretching, asymmetrical | [87, 88] | |
| 3087 | C-N stretching + overtone of N-H in-plane bending | [86, 87, 89, 88] | |
| 3300 | N-H stretching | [86, 87, 84, 88, 90] |